# Observer-robust energy condition verification for warp drive spacetimes

### v1

@@ 4.2 WarpFactory comparison
i.e. the fraction of all grid points where the Eulerian margin is non-negative (condition appears satisfied) but the robust ground truth is negative (condition violated). For Type I points (the dominant case; see Table 8), the ground truth $m_{\text{rob}}$ is the algebraic slack (Section 2.2), which is cap-independent; consequently, $f_{\text{miss}}$ is insensitive to $\zeta_{\max}$ in practice. Both $f_{\text{miss}}$ and “% violated” are domain-restricted diagnostics: their values depend on the sampled region and grid density, and are primarily meaningful for comparing the Eulerian and robust analyses within the same domain (not as intrinsic spacetime invariants). Table 3 summarizes $f_{\text{miss}}$ for all metrics at $v_{s}=0.5$. The Rodal metric shows substantial missed violations across NEC, WEC, SEC, and DEC. SEC misses are notably widespread: Van Den Broeck has $1.2\%$ SEC missed despite nonzero NEC/WEC/DEC misses, demonstrating that SEC violations are genuinely independent of the other conditions. This independence arises because SEC involves the trace $\rho+\sum_{i}p_{i}$, which can be negative even when each individual $\rho+p_{i}\geq 0$ (NEC satisfied). The violation severity can differ substantially between Eulerian and optimized observers even when the violation sets coincide. Table values are percentages over the full $50^{3}$ volume rounded to $0.1\%$; slice plots may show a few points even when the rounded volume fraction is $0.0\%$. In unrounded values, WarpShell shows NEC total violated at $0.259\%$ and WEC total violated at $0.270\%$. A slight WEC >> NEC difference is expected because the WEC is strictly stronger than the NEC: WEC satisfaction implies NEC satisfaction, but WEC can be violated even when NEC holds. At this sub-percent level, exact counts can also be influenced by finite-precision effects near the violation boundary and by the fact that our non-Type I diagnostics are $\zeta_{\max}$-capped. At the 0.1% reporting precision of Tables 3 and 4, both values round to the same figure; the underlying sub-percent difference is visible only at higher precision. Table 4 shows the velocity dependence for the two metrics with the largest discrepancies.
@@
Table 3: Energy condition analysis at $v_{s}=0.5$ on a $50^{3}$ grid ($\zeta_{\max}=5$ throughout). Total: percentage of grid points violating each condition (hybrid margin: algebraic slacks at Type I points, $\zeta_{\max}$-capped extrema at non-Type I). Missed: percentage of grid points where the Eulerian analysis fails to detect a violation found by the hybrid check. The Eulerian violated fraction equals the total violated fraction minus the missed fraction. A missed fraction of $0.0\%$ does not imply zero violations; it means all violations are detected by the Eulerian frame. Bold entries highlight missed fractions $\geq 1\%$. Lentz and WarpShell are under-resolved ($\ll\!1$ and $\sim\!0.3$ cells across the wall, respectively; Table 2); their total and missed fractions should be interpreted as lower bounds.
@@ 4.3 Velocity dependence
For the Alcubierre metric (Figure 7a), the Eulerian frame detects all NEC and WEC violation points at every velocity tested (missed $0.0\%$). The robust NEC violation fraction grows from 1.73% at $v_{s}=0.1$ to 2.21% at $v_{s}=0.99$, and the robust WEC violation fraction tracks NEC identically; the violation regions coincide with the Eulerian analysis, but the observer-optimized margins are consistently more negative, indicating greater violation severity for boosted observers.
For the Rodal metric, the discrepancy is far more dramatic. The missed DEC violation fraction is approximately 28.5% at all velocities, indicating that the Eulerian DEC analysis fails to detect violations at roughly 28.5% of all grid points. Missed WEC violations reach  ${\sim}\,15\%$ across the velocity sweep, growing from 14.9% at $v_{s}=0.1$ to 15.6% at $v_{s}=0.99$. Even the NEC missed fraction is non-zero for Rodal, ranging from 1.26% to 1.61%. The algebraic DEC slacks at missed points are shallow: the observer-optimized minimum is $2.0\times 10^{-6}$ (Section 8), and deterministic Fibonacci sampling saturates at $\sim\!93$% detection even at $10^{4}$ directions per point (Figure 13), confirming that the undetected 28.5% comprises violations near the satisfaction boundary rather than deeply negative points.
@@ 6 Results: Kinematic scalars
Figure 10 shows the kinematic scalars for the Alcubierre metric at $v_{s}=0.5$. The expansion $\theta$ has the characteristic bipolar structure: positive expansion ahead of the bubble (space expanding) and negative expansion behind (space contracting), with $\theta=0$ inside the flat region. The shear is concentrated at the bubble wall and peaks where the expansion gradient is largest. These kinematic features are closely connected to the exotic matter requirements. For a geodesic, irrotational timelike congruence, the Raychaudhuri equation $\dot{\theta}=-\tfrac{1}{3}\theta^{2}-\sigma_{ab}\sigma^{ab}-R_{ab}\,u^{a}u^{b}$ links the expansion evolution to $R_{ab}\,u^{a}u^{b}$, and hence (via the Einstein equations) to the SEC. The Eulerian congruence is in general accelerated (with acceleration $a^{b}=\nabla_{b}\ln\alpha$), so the connection is qualitative rather than a strict implication; a similar link to the NEC holds for null geodesic congruences.
@@ 7 Resolution stability
Table 2 characterizes the wall resolution for each metric at $50^{3}$. Rodal is the only metric that comfortably exceeds the 4-cell resolution threshold ($\sim\!6$ cells); Alcubierre, VdB, and Natário are marginally resolved ($\sim\!1.4$ cells), Lentz is severely under-resolved ($\sim\!0.02$ cells), and WarpShell concentrates its thin-shell transition within $\lesssim\!1$ cell. Autodiff computes exact curvature at each grid point regardless of wall resolution. To verify that the headline missed-violation fractions are not resolution artifacts, we run the same three-resolution sweep ($25^{3}$$50^{3}$$100^{3}$) for the Rodal ($R=100$, $\sigma=0.03$) and WarpShell ($R_{1}=0.5$, $R_{2}=1.0$) metrics. At $100^{3}$ we use Eulerian-only evaluation due to the prohibitive cost of full observer optimization at $10^{6}$ points. The minimum NEC margin converges at second order for Rodal: extrapolation gives $2.58\times 10^{-6}$ (503 value: $2.62\times 10^{-6}$, within 1.5%). For WarpShell, the Eulerian minimum NEC margin increases in absolute magnitude with resolution ($2.5\times 10^{25}$ at $25^{3}$, $2.3\times 10^{33}$ at $50^{3}$, $2.1\times 10^{34}$ at $100^{3}$), reflecting the shell boundary extremes that are better resolved at finer grids; Richardson extrapolation yields a negative convergence order ($p\approx-3.0$), so we report per-resolution values rather than an extrapolated limit. The Rodal integrated violation does not converge monotonically, consistent with the broad Rodal wall ($\sigma=0.03$, 10–90% transition width $73.24$ coordinate units spanning $\sim\!6.0$ grid cells; Table 2), which is marginally resolved at the $50^{3}$ grid but whose small gradients require higher resolution to stabilize extensive (volume-integrated) quantities.
@@ 7.1 Curvature identity validation
As an independent check of the autodiff curvature chain, we verify that the computed Riemann tensor satisfies fundamental geometric identities at representative points for the Schwarzschild and Alcubierre metrics. The three algebraic Riemann symmetries (antisymmetry in the first and second index pairs, $R_{abcd}=-R_{bacd}$ and $R_{abcd}=-R_{abdc}$, and pair interchange symmetry, $R_{abcd}=R_{cdab}$) hold with maximum residual $10^{-12}$ across all test points. The first Bianchi identity ($R_{abcd}+R_{acdb}+R_{adbc}=0$) is satisfied to the same precision. The contracted Bianchi identity ($\nabla_{a}G^{ab}=0$) is verified to a residual $10^{-6}$ (absolute, normalized by $\|G^{ab}\|_{\max}$ this corresponds to a relative residual $\lesssim 10^{-6}$) via autodiff evaluation of the covariant divergence; the larger residual ($10^{-6}$ vs. $10^{-12}$ for the algebraic identities) is expected because this identity involves third derivatives of the metric, where floating-point cancellation is more severe. All computations use 64-bit floating point (jax.config.update("jax_enable_x64", True)). Since $T_{ab}=G_{ab}/(8\pi)$ and $G_{ab}$ inherits the symmetry of $R_{ab}$ and $g_{ab}$, the stress-energy tensor is symmetric by construction; spot-checked residuals $|T_{ab}-T_{ba}|_{\max}<10^{-14}$ confirm this numerically. These results confirm that the forward-mode autodiff chain correctly preserves the geometric structure of the curvature tensor.
@@ 7.3 Classification breakdown and conditional miss rate
Table 8 reports the Hawking–Ellis type breakdown for all seven metrics at $v_{s}=0.5$ (Schwarzschild at $v_{s}=0$). We employ a two-tier imaginary tolerance (Section 8): (i) absolute, $|\mathrm{Im}\,\lambda_{i}|<\tau\cdot s$ with $s=\max(|\mathrm{Re}\,\lambda|,1)$, and (ii) relative, $|\mathrm{Im}\,\lambda_{i}|<3\times 10^{-3}\cdot\max|\mathrm{Re}\,\lambda|$ (unclamped), which catches split degenerate pairs at large $\|T\|$. Rodal, Schwarzschild, and Lentz are $\geq\!99.9$% Type I. Alcubierre, Natário, and VdB are $\sim\!97$9898% Type I, with the remaining points classified as Type IV at the transition zone between the warp bubble and flat space. WarpShell achieves $\sim\!99.6$% Type I, with $\sim\!0.3$% Type II at the shell boundary and $\sim\!0.1$% Type IV. WarpShell serves primarily as a numerical regularity stress test: its prescribed $C^{2}$ transition functions generate extreme curvature scales that exercise the classifier, optimizer, and convergence diagnostics in a worst-case regime; the resulting EC margins should be understood as properties of the regularized implementation, not as predictions for an idealized thin-shell spacetime. The residual non-Type I WarpShell points are concentrated in the shell transition region where $\|T^{a}{}_{b}\|$ reaches $\mathcal{O}(10^{25})$ (at $50^{3}$ resolution; the eigenvalue scale grows with resolution as the grid resolves the thin shell). A dedicated tetrad-normalization diagnostic confirms that the observer-orthonormal tetrad is well-normalized ($\max|g(k,k)|<10^{-6}$ and $\max|\eta_{IJ}-g(e_{I},e_{J})|<10^{-6}$ at >99.8>99.8% of grid points), and that NEC contractions $T_{ab}k^{a}k^{b}$ are the same order of magnitude as the eigenvalues of $T^{a}{}_{b}$ (ratio $\sim 0.4$). At finer grids the eigenvalue scale itself increases (reaching $\sim\!10^{33}$ at $100^{3}$; Section 7), and observer boosting at $\zeta_{\max}=5$ ($\gamma\approx 74$) further amplifies contractions by $\cosh^{2}\!\zeta\approx 5500$, bridging the gap between the $50^{3}$ eigenvalue scale and the reported $\mathcal{O}(10^{30})$$\mathcal{O}(10^{35})$ margins. These therefore reflect genuine curvature amplitudes at thin-shell grid points, not a normalization artefact. At these points, eigenvalue pairs of the non-symmetric mixed tensor split as $\lambda\pm\varepsilon i$ with $|\varepsilon|/|\lambda|\sim 10^{-3}$$10^{-2}$. The two-tier tolerance (Section 8) recovers the majority of such points as effectively real; the residual 0.10.1% Type IV fraction is comparable to that of the other warp metrics. The classifier is conservative by design: a point is labeled Type I only when eigenvalues pass either tier of the imaginary-part check. Any ambiguous point defaults to the optimizer pathway. Misclassifying Type I as non-Type I incurs only a performance penalty (unnecessary optimization); the reverse (labeling non-Type I as Type I) would require complex eigenvalues to simultaneously satisfy both the absolute and relative tolerance, which the tolerance sensitivity analysis (Section 8) confirms is extremely unlikely at the tested thresholds. At non-Type I points the capped extrema are used; at Type I points the algebraic slacks dominate.
@@
This quantity measures how “leaky” the Eulerian frame is at points where violations actually exist, complementing the unconditional $f_{\text{miss}}$ which is diluted by the (potentially large) number of satisfied points.

### v2

@@ 4.2 WarpFactory comparison
i.e. the fraction of all grid points where the Eulerian margin is non-negative (condition appears satisfied) but the robust ground truth is negative (condition violated). For Type I points (the dominant case; see Table 8), the ground truth $m_{\text{rob}}$ is the algebraic slack (Section 2.2), which is cap-independent; consequently, $f_{\text{miss}}$ is insensitive to $\zeta_{\max}$ in practice. Both $f_{\text{miss}}$ and “% violated” are domain-restricted diagnostics: their values depend on the sampled region, grid density, and—crucially—the wall thickness of each metric. Metrics with broader walls (e.g. Rodal, $\sigma=0.03$, wall width $\approx 73$ cells; Table 2) naturally occupy more grid volume, inflating volume-based violation fractions relative to thin-wall metrics. These fractions are therefore not directly comparable across metrics as a severity measure; they are primarily meaningful for comparing the Eulerian and robust analyses within the same domain and metric. The conditional miss rate $f_{\text{miss}|\text{viol}}\equiv(\text{missed})/(\text{total violated})$, which normalizes by the number of actually-violated points, is the more appropriate cross-metric diagnostic for our central claim. Table 3 summarizes $f_{\text{miss}}$ for all metrics at $v_{s}=0.5$. The Rodal metric shows substantial missed violations across NEC, WEC, SEC, and DEC. SEC misses are notably widespread: Van Den Broeck has $1.2\%$ SEC missed despite nonzero NEC/WEC/DEC misses, demonstrating that SEC violations are genuinely independent of the other conditions. This independence arises because SEC involves the trace $\rho+\sum_{i}p_{i}$, which can be negative even when each individual $\rho+p_{i}\geq 0$ (NEC satisfied). The violation severity can differ substantially between Eulerian and optimized observers even when the violation sets coincide. Table values are percentages over the full $50^{3}$ volume rounded to $0.1\%$; slice plots may show a few points even when the rounded volume fraction is $0.0\%$. In unrounded values, WarpShell shows NEC total violated at $0.259\%$ and WEC total violated at $0.270\%$. A slight WEC >> NEC difference is expected because the WEC is strictly stronger than the NEC: WEC satisfaction implies NEC satisfaction, but WEC can be violated even when NEC holds. At this sub-percent level, exact counts can also be influenced by finite-precision effects near the violation boundary and by the fact that our non-Type I diagnostics are $\zeta_{\max}$-capped. At the 0.1% reporting precision of Tables 3 and 4, both values round to the same figure; the underlying sub-percent difference is visible only at higher precision. Table 4 shows the velocity dependence for the two metrics with the largest discrepancies.
@@
Table 3: Energy condition analysis at $v_{s}=0.5$ on a $50^{3}$ grid ($\zeta_{\max}=5$ throughout). Total: percentage of grid points violating each condition (hybrid margin: algebraic slacks at Type I points, $\zeta_{\max}$-capped extrema at non-Type I). Missed: percentage of grid points where the Eulerian analysis fails to detect a violation found by the hybrid check. The Eulerian violated fraction equals the total violated fraction minus the missed fraction. A missed fraction of $0.0\%$ does not imply zero violations; it means all violations are detected by the Eulerian frame. Bold entries highlight missed fractions $\geq 1\%$. Volume fractions scale with wall thickness (Table 2); cross-metric comparisons should use the conditional miss rate $f_{\text{miss}|\text{viol}}$ (missed/total violated). Lentz and WarpShell are under-resolved ($\ll\!1$ and $\sim\!0.3$ cells across the wall, respectively; Table 2); their total and missed fractions should be interpreted as lower bounds.
@@ 4.3 Velocity dependence
For the Alcubierre metric (Figure 7a), the Eulerian frame detects all NEC and WEC violation points at every velocity tested (missed $0.0\%$). The robust NEC violation fraction grows from 1.73% at $v_{s}=0.1$ to 2.21% at $v_{s}=0.99$, and the robust WEC violation fraction tracks NEC identically; the violation regions coincide with the Eulerian analysis, but the hybrid margins are consistently more negative, indicating greater violation severity for boosted observers.
For the Rodal metric, the discrepancy is far more dramatic. The missed DEC violation fraction is approximately 28.5% at all velocities, indicating that the Eulerian DEC analysis fails to detect violations at roughly 28.5% of all grid points. Missed WEC violations reach  ${\sim}\,15\%$ across the velocity sweep, growing from 14.9% at $v_{s}=0.1$ to 15.6% at $v_{s}=0.99$. Even the NEC missed fraction is non-zero for Rodal, ranging from 1.26% to 1.61%. The algebraic DEC slacks at missed points are shallow: the capped extremum minimum is $2.0\times 10^{-6}$ (Section 8), and deterministic Fibonacci sampling saturates at $\sim\!93$% detection even at $10^{4}$ directions per point (Figure 13), confirming that the undetected 28.5% comprises violations near the satisfaction boundary rather than deeply negative points.
@@ 6 Results: Kinematic scalars
Figure 10 shows the kinematic scalars for the Alcubierre metric at $v_{s}=0.5$. The expansion $\theta$ has the characteristic bipolar structure: positive expansion ahead of the bubble (space expanding) and negative expansion behind (space contracting), with $\theta=0$ inside the flat region. The shear is concentrated at the bubble wall and peaks where the expansion gradient is largest. These kinematic features are closely connected to the exotic matter requirements. For a geodesic, irrotational timelike congruence, the Raychaudhuri equation $\dot{\theta}=-\tfrac{1}{3}\theta^{2}-\sigma_{ab}\sigma^{ab}-R_{ab}\,u^{a}u^{b}$ links the expansion evolution to $R_{ab}\,u^{a}u^{b}$, and hence (via the Einstein equations) to the SEC. For the unit-lapse warp metrics in our suite (Alcubierre, Lentz, Natário, Van Den Broeck, Rodal), $\alpha\equiv 1$ implies $a^{b}=\nabla^{b}\ln\alpha=0$, so the Eulerian congruence is geodesic and this link is exact. For WarpShell ($\alpha\neq 1$) the congruence is accelerated and the connection is qualitative rather than a strict implication; a similar link to the NEC holds for null geodesic congruences.
@@ 7 Resolution stability
Table 2 characterizes the wall resolution for each metric at $50^{3}$. Rodal is the only metric that comfortably exceeds the 4-cell resolution threshold ($\sim\!6$ cells); Alcubierre, VdB, and Natário are marginally resolved ($\sim\!1.4$ cells), Lentz is severely under-resolved ($\sim\!0.02$ cells), and WarpShell concentrates its thin-shell transition within $\lesssim\!1$ cell. Autodiff computes exact curvature at each grid point regardless of wall resolution. To verify that the headline missed-violation fractions are not resolution artifacts, we run the same three-resolution sweep ($25^{3}$$50^{3}$$100^{3}$) for the Rodal ($R=100$, $\sigma=0.03$) and WarpShell ($R_{1}=0.5$, $R_{2}=1.0$) metrics. At $100^{3}$ we use Eulerian-only evaluation due to the prohibitive cost of full observer optimization at $10^{6}$ points. The minimum NEC margin converges at second order for Rodal: extrapolation gives $2.58\times 10^{-6}$ ($50^{3}$ value: $2.62\times 10^{-6}$, within 1.5%). For WarpShell, the Eulerian minimum NEC margin increases in absolute magnitude with resolution ($2.5\times 10^{25}$ at $25^{3}$, $2.3\times 10^{33}$ at $50^{3}$, $2.1\times 10^{34}$ at $100^{3}$), reflecting the shell boundary extremes that are better resolved at finer grids; Richardson extrapolation yields a negative convergence order ($p\approx-3.0$), so we report per-resolution values rather than an extrapolated limit. The Rodal integrated violation does not converge monotonically, consistent with the broad Rodal wall ($\sigma=0.03$, 10–90% transition width $73.24$ coordinate units spanning $\sim\!6.0$ grid cells; Table 2), which is marginally resolved at the $50^{3}$ grid but whose small gradients require higher resolution to stabilize extensive (volume-integrated) quantities.
@@ 7.1 Curvature identity validation
As an independent check of the autodiff curvature chain, we verify that the computed Riemann tensor satisfies fundamental geometric identities at representative points for the Schwarzschild and Alcubierre metrics. The three algebraic Riemann symmetries (antisymmetry in the first and second index pairs, $R_{abcd}=-R_{bacd}$ and $R_{abcd}=-R_{abdc}$, and pair interchange symmetry, $R_{abcd}=R_{cdab}$) hold with maximum residual $10^{-12}$ across all test points. The first Bianchi identity ($R_{abcd}+R_{acdb}+R_{adbc}=0$) is satisfied to the same precision. The contracted Bianchi identity ($\nabla_{a}G^{ab}=0$) is verified to a residual $10^{-6}$ (absolute, normalized by $\|G^{ab}\|_{\max}$; this corresponds to a relative residual $\lesssim 10^{-6}$) via autodiff evaluation of the covariant divergence; the larger residual ($10^{-6}$ vs. $10^{-12}$ for the algebraic identities) is expected because this identity involves third derivatives of the metric, where floating-point cancellation is more severe. All computations use 64-bit floating point (jax.config.update("jax_enable_x64", True)). Since $T_{ab}=G_{ab}/(8\pi)$ and $G_{ab}$ inherits the symmetry of $R_{ab}$ and $g_{ab}$, the stress-energy tensor is symmetric by construction; spot-checked residuals $|T_{ab}-T_{ba}|_{\max}<10^{-14}$ confirm this numerically. These results confirm that the forward-mode autodiff chain correctly preserves the geometric structure of the curvature tensor.
@@ 7.3 Classification breakdown and conditional miss rate
Table 8 reports the Hawking–Ellis type breakdown for all seven metrics at $v_{s}=0.5$ (Schwarzschild at $v_{s}=0$). We employ a two-tier imaginary tolerance (Section 8): (i) absolute, $|\mathrm{Im}\,\lambda_{i}|<\tau\cdot s$ with $s=\max(|\mathrm{Re}\,\lambda|,1)$, and (ii) relative, $|\mathrm{Im}\,\lambda_{i}|<3\times 10^{-3}\cdot\max|\mathrm{Re}\,\lambda|$ (unclamped), which catches split degenerate pairs at large $\|T\|$. Rodal, Schwarzschild, and Lentz are $\geq\!99.9$% Type I. Alcubierre, Natário, and VdB are $\sim\!97$9898% Type I, with the remaining points classified as Type IV at the transition zone between the warp bubble and flat space. WarpShell achieves $\sim\!99.6$% Type I, with $\sim\!0.3$% Type II at the shell boundary and $\sim\!0.1$% Type IV. WarpShell serves primarily as a numerical regularity stress test: its prescribed $C^{2}$ transition functions generate extreme curvature scales that exercise the classifier, optimizer, and convergence diagnostics in a worst-case regime; the resulting EC margins should be understood as properties of the regularized implementation, not as predictions for an idealized thin-shell spacetime. The residual non-Type I WarpShell points are concentrated in the shell transition region where $\|T^{a}{}_{b}\|$ reaches $\mathcal{O}(10^{25})$ (at $50^{3}$ resolution; the eigenvalue scale grows with resolution as the grid resolves the thin shell). A dedicated tetrad-normalization diagnostic confirms that the observer-orthonormal tetrad is well-normalized ($\max|g(k,k)|<10^{-6}$ and $\max|\eta_{IJ}-g(e_{I},e_{J})|<10^{-6}$ at >99.8>99.8% of grid points), and that NEC contractions $T_{ab}k^{a}k^{b}$ are the same order of magnitude as the eigenvalues of $T^{a}{}_{b}$ (ratio $\sim 0.4$). At finer grids the eigenvalue scale itself increases (reaching $\sim\!10^{33}$ at $100^{3}$; Section 7), and observer boosting at $\zeta_{\max}=5$ ($\gamma\approx 74$) further amplifies contractions by $\cosh^{2}\!\zeta\approx 5500$, bridging the gap between the $50^{3}$ eigenvalue scale and the reported $\mathcal{O}(10^{30})$$\mathcal{O}(10^{35})$ margins. These therefore reflect genuine curvature amplitudes at thin-shell grid points, not a normalization artifact. At these points, eigenvalue pairs of the non-symmetric mixed tensor split as $\lambda\pm\varepsilon i$ with $|\varepsilon|/|\lambda|\sim 10^{-3}$$10^{-2}$. The two-tier tolerance (Section 8) recovers the majority of such points as effectively real; the residual 0.10.1% Type IV fraction is comparable to that of the other warp metrics. The classifier is conservative by design: a point is labeled Type I only when eigenvalues pass either tier of the imaginary-part check. Any ambiguous point defaults to the optimizer pathway. Misclassifying Type I as non-Type I incurs only a performance penalty (unnecessary optimization); the reverse (labeling non-Type I as Type I) would require complex eigenvalues to simultaneously satisfy both the absolute and relative tolerance, which the tolerance sensitivity analysis (Section 8) confirms is extremely unlikely at the tested thresholds. At non-Type I points the capped extrema are used; at Type I points the algebraic slacks dominate.
@@
This quantity measures how “leaky” the Eulerian frame is at points where violations actually exist, complementing the unconditional $f_{\text{miss}}$, which is diluted by the (potentially large) number of satisfied points.